OpenStudy (anonymous):

What is the solution set of 6x 2 - 24 = 0? {-2, 2} {-2} {2}

OpenStudy (jit4won):

{-2, 2}

OpenStudy (lgbasallote):

factor out 6 then you'll have a difference of two squares

OpenStudy (lgbasallote):

@jit4won we are not allowed to give direct answers here. we try hard to teach our askers :) i hope you learn better ^_^

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 6(x^2 - 4) = 0 }\) \(\Large \color{purple}{\rightarrow x^2 - 4 = 0 }\) \(\Large \color{purple}{\rightarrow x^{2} = 4 }\) \(\Large \color{purple}{\rightarrow x = \sqrt{4} }\) \(\Large \color{purple}{\rightarrow x = 2(or -2) }\)

OpenStudy (lgbasallote):

that goes for you as well @ParthKohli i would suggest giving a solution but NOT an answer :) ~Thanks

Parth (parthkohli):

@lgbasallote When I sent him the code of conduct, he said I know what to do and he's continuing the direct answer thingy